# Leveraging Temporal Contextualization for Video Action Recognition

### v1

@@ 3.3 Video-conditional Prompting (VP)
We present the Video-conditional Prompting (VP) module, which aims to further leverage the precisely aggregated temporal information in visual domain for text encoding. We impose a cross-attention between prompt vectors and context tokens to enrich the information in the prompt vectors, and the detailed procedure is illustrated in Fig. 5. Let $\textbf{c}^{l-1}$ and $\textbf{p}^{l-1}$ be class name tokens and learnable prompt vectors from the $(l-1)^{\text{th}}$ layer of the text encoder, respectively. We derive temporally contextualized prompt vectors $\hat{\mathbf{p}}^{l-1}$ by passing the layer-normalized prompt tokens and context tokens through a cross-attention layer as follows:
Figure 5: Video-conditional Prompting (VP) module. Video information from the context tokens is injected to the text prompt vectors based on a cross-attention mechanism, generating instance-level prompts that support the lack of textual semantics.
@@
$\hat{\mathbf{p}}^{l-1}$ $=\text{MHCA}(\text{LN}_{p}(\mathbf{p}^{l-1}),\text{LN}_{s}(\mathbf{s}_{\text{proj}}^{l}))+\mathbf{p}^{l-1}$ (15)
where $\text{SG}(\cdot)$ is a stop-gradient function, $\mathbf{W}_{\text{vis}}$ is a weight matrix of CLIP to linearly project vision representations to a common vision-language latent space, and $\text{MHCA}(\cdot,\cdot)$ is a multi-head cross-attention opearation for interactions across modalities, accepting text prompt vectors as queries and vision features as keys and values. Finally, the contextualized prompt vector $\hat{\mathbf{p}}^{l-1}$ are updated by FFN:
$\tilde{\mathbf{p}}^{l-1}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{p}}^{l-1})+\hat{\mathbf{p}}^{l-1}$ (16)
The VP module $f_{\theta_{\text{VP}}}(\cdot,\cdot)$ is defined by a composition of Eq. (15) and Eq. (16), and executed before the last layer of text encoder $f_{\theta_{c}}$. Finally, the new formulation of our encoding process in the text modality is given by
@@
The resulting instance-conditioned prompt vectors effectively support the lack of textual semantics in action recognition datasets. The impact of our VP design is verified in the experiments in Sec. 5.
@@ 3.4 Training Objective
TC-CLIP learns to maximize the similarity of video representations $\mathbf{v}$ and text representations $\mathbf{c}$ for matching pairs in a mini-batch via the cross-entropy loss:
$\mathcal{L}=-\sum_{i}\log\frac{\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{j})/\tau)},$ (18)
where $\tau$ is a learnable temperature parameter. Our model is fully fine-tuned in an end-to-end manner.
@@ 4 Experiments
4.1 Experimental Setup
Datasets. We conduct experiments over 5 video benchmarks: Kinetics-400 [16] & 600 [2], HMDB-51 [18], UCF-101 [28], and Something-Something v2 (SSv2) [8]. Following [26], our evaluation protocol includes zero-shot, few-shot, base-to-novel generalization, and fully-supervised action recognition.
Implementation details. We adopt CLIP with ViT-B/16 for all experiments. Our baseline is ViFi-CLIP [26]. During the training, we sample 16 frames to form a video clip. During the evaluation, two temporal clips with one spatial crop (2 $\times$ 1 view) per video are sampled to produce a prediction unless otherwise stated. For training recipes, we follow [26] for zero-shot, few-shot, and fully-supervised settings and follow [9] for base-to-novel generalization. Following [33, 20, 5, 9], we adopt linear weight-space ensembling of our fine-tuned model and CLIP for the zero-shot protocol. We set the number of merged tokens $r$ in each bipartite matching iteration to 100, and the weight of the VP module is initialized with the weight from the last layer of the CLIP text encoder. All models are trained using 4 NVIDIA Tesla V100 GPUs. See supplementary material for more details.
Table 1: Comparison with state-of-the-arts on zero-shot action recognition. All the models are trained on Kinetics-400 and directly evaluated on other datasets. † denotes the methods with weight-space ensemble. Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
Method HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
@@ 4.1 Experimental Setup
ViFi-CLIP [26] 51.3 $\pm$ 0.6 76.8 $\pm$ 0.7 71.2 $\pm$ 1.0 92.2 $\pm$ 0.3 66.4
Open-VCLIP [33] † 53.9 $\pm$ 1.2 83.4 $\pm$ 1.2 73.0 $\pm$ 0.8 93.2 $\pm$ 0.1 70.1
TC-CLIP (Ours) † 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 94.4 $\pm$ 0.2 71.0
@@
MAXI [20] † 52.3 $\pm$ 0.7 78.2 $\pm$ 0.8 71.5 $\pm$ 0.8 92.5 $\pm$ 0.4 67.3
OST [5] † 55.9 $\pm$ 1.2 79.7 $\pm$ 1.1 75.1 $\pm$ 0.6 94.6 $\pm$ 0.2 70.2
FROSTER [9] † 54.8 $\pm$ 1.3 84.8 $\pm$ 1.1 74.8 $\pm$ 0.9 - 71.5
TC-CLIP (Ours) † 56.0 $\pm$ 0.3 85.4 $\pm$ 0.8 78.1 $\pm$ 1.0 95.7 $\pm$ 0.3 73.2
4.2 Quantitative Comparison
We mainly compare our method with CLIP-based video recognition models: Vanilla CLIP [24], ActionCLIP [30], A5 [13], X-CLIP [22], Vita-CLIP [32], ViFi-CLIP [26], Open-VCLIP [33], OST [5], and FROSTER [9]. To fairly compare our method with approaches that utilize Large Language Model (LLM)-based text augmentation [20, 5, 9], we divide our results into two versions: one using the original action category names (colored with blue) and the other adopting the LLM-rephrased category names from FROSTER [9] (colored with purple). Note that experiments on the SSv2 dataset do not involve LLM-rephrasing.
[…]
@@ 4.2 Quantitative Comparison
Table 3: Comparison with state-of-the-arts on base-to-novel generalization. All the models are directly fine-tuned from CLIP. † denotes the results are taken from [9]. Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
@@
Few-shot action recognition. We demonstrate the learning capacity of our method under a challenging few-shot scenario. In this setting, models are directly fine-tuned from CLIP on HMDB-51, UCF-101, and SSv2 using $K$-shot samples, where $K$ is set to 2, 4, 8, and 16. Table 2 shows the comparison with state-of-the-art video action recognition methods on the few-shot setting. Our method achieves state-of-the-art performance with an average accuracy of 54.8%, surpassing our baseline model ViFi-CLIP [26] by 1.9%p.
Base-to-novel generalization. Similar to the few-shot setting, models are directly fine-tuned from CLIP on the base classes of K-400, HMDB-51, UCF-101, and SSv2, and evaluated for both base and novel classes. The top-1 accuracy on base and novel classes, and their harmonic mean (HM) are reported. Table 3 shows that our model achieves state-of-the-art performance on the novel classes and HM across all datasets. In particular, TC-CLIP achieves an average HM accuracy of 58.5%, surpassing previous state-of-the-arts by 2.0%p.
@@
Table 4: Comparison with state-of-the-arts on fully-supervised action recognition on Kinetics-400. Views are denoted with temporal clips $\times$ spatial crops. † means that computational costs are measured in our environment using a single V100 GPU. The results of other methods are taken from [26].
Method Top-1 Top-5 Frames Views GFLOPs Params (M)
ActionCLIP [30] 83.8 96.2 32 10 $\times$ 3 563 168.5
X-CLIP [22] 84.7 96.8 16 4 $\times$ 3 287 131.5
Vita-CLIP [22] † 82.9 96.3 16 4 $\times$ 3 307 185.3
ViFi-CLIP [26] 83.9 96.3 16 4 $\times$ 3 281 124.7
OST [5] 83.2 - 16 1 $\times$ 1 - -
TC-CLIP (Ours) 85.2 96.9 16 4 $\times$ 3 304 127.5
@@
Fully-supervised action recognition. Table 4 shows performance comparison results under fully-supervised setting, where the models are both trained and evaluated on the K-400 dataset. Our model achieves top-1 accuracy of 85.2% in the K-400 validation split, improving 1.3%p over our baseline ViFi-CLIP [26].
@@
Table 5: Component-wise ablation on the zero-shot setting. ($\Delta$) denotes the average top-1 accuracy improvement over baseline.
Case TC VP WE HMDB-51 UCF-101 K-600 All ($\Delta$)
Baseline 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 67.3
(a) w/o Video-conditional Prompting (VP) 54.3 $\pm$ 0.6 81.9 $\pm$ 1.0 75.5 $\pm$ 0.5 70.6 ($+$3.3)
(b) w/o Temporal Contextualization (TC) 53.2 $\pm$ 0.8 82.0 $\pm$ 0.9 74.7 $\pm$ 0.7 70.0 ($+$2.7)
(c) w/o Weight-space Ensembling (WE) 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 68.9 ($+$1.6)
TC-CLIP 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 71.0 ($+$3.7)
Table 6: Temporal Contextualization (TC) ablation on the few-shot setting. SE, JE, and Bias denote spatial embedding, joint space-time embedding, and bias (Eq. 10). Default settings are marked in gray.
(a) Seed token selection metric.
@@
No selection 62.8 89.8 9.7 54.1
Head-wise key norm 62.3 89.8 9.8 54.0
Averaged key norm 62.5 89.4 9.3 53.7
Head-wise CLS attention 63.4 89.9 9.7 54.3
Averaged CLS attention 63.4 90.2 9.9 54.5
(b) Token aggregation method.
Case HMDB UCF SSv2 All
No merge 57.2 85.6 7.7 50.2
Temporal tube 57.9 86.4 8.3 50.9
Random merge 58.8 87.1 7.5 51.2
K-means 61.7 89.7 9.0 53.4
Bipartite matching 63.4 90.2 9.9 54.5
(c) Seed token ratio $\alpha$.
@@
256 60.9 89.5 9.3 53.2
(e) Positional embedding.
Case HMDB UCF SSv2 All
SE 62.9 90.0 9.8 54.2
JE 63.2 90.2 9.8 54.4
SE+Bias 63.4 90.2 9.9 54.5
JE+Bias 62.9 90.2 9.8 54.3
5 Analysis and Discussion
This section examines the design choices and impact of each component in our model: Temporal Contextualization (TC) and Video-conditional Prompting (VP). We mainly adopt the zero- and few-shot settings and report the average of top-1 accuracy from $K=2,4,8,16$ for the $K$-shot setup.
5.1 Component-wise Ablation
Table 5 shows the impact of TC, VP, and weight-space ensembling (WE) on our baseline in the zero-shot setting. (a) Adopting TC with WE leads to an average gain of 3.3%p, while (b) incorporating VP with WE results in a 2.7%p improvement, highlighting the significance of TC in our framework. (c) When both VP and TC are applied to the baseline, we observed a 1.6%p improvement on average. Finally, by integrating all three components, TC-CLIP achieves a total 3.7%p gain on average.
5.2 Temporal Contextualization (TC) Ablation
Table 6 analyzes the design choice of TC. Experiments are conducted on the few-shot setting using the baseline model with TC.
Seed token selection metric. Using [CLS] attention score averaged from multiple heads is the best choice, resulting in a 0.4%p gain on average compared to the no-selection case, i.e., using all tokens for aggregation.
Token aggregation methods. Using seed tokens without merging leads to inferior performance, encountering the extrapolation issue. When comparing naive aggregation methods, random merge surpasses temporal tube since random merge has more chance to aggregate relevant tokens while temporal tube are enforced to group tokens of the same spatial positions. Finally, we compared the clustering algorithms and selected bipartite matching as our default choice.
Seed token selection ratio $\alpha$. While TC is not sensitive to the choice of $\alpha$, we picked $\alpha=0.3$ as our default value, i.e., using 30% of total tokens.
Context token number $k$. By varying the value of $k$, we observed that selecting a modest amount of merging degree is important for achieving the best performance. An excessively low $k$ leads to the lost of spatial details like part-wise information, while an overly high $k$ results in inadequate contextualization of tokens into relevant sets. Therefore, we choose $k=96$ as our default setting.
Positional embedding. Among spatial embedding (SE), joint space-time embedding (JE), and our proposed learnable bias (Eq. 10), we observe that using SE with bias yields the best result. We conjecture that the bias effectively consolidates the local frame-level information and global video-level information in a layer-wise and head-wise manner.
Table 7: Video-conditional Prompting (VP) ablation. Default settings are marked in gray.
(a) Vision token selection (few-shot).
Case HMDB UCF SSv2 All
CLS tokens 63.3 90.3 9.8 54.4
GAP tokens 63.7 90.5 9.8 54.7
Context tokens 63.9 90.7 9.8 54.8
(b) Number of prompt vectors (few-shot).
Case HMDB UCF SSv2 All
2 62.0 90.1 9.4 53.8
4 63.9 90.7 9.8 54.8
8 63.7 90.4 9.8 54.7
(c) Text prompting design (zero-shot). All the models are evaluated without weight ensembling.
@@ 5.2 Temporal Contextualization (TC) Ablation
5.3 Video-conditional Prompting (VP) Ablation
In Table 7, we analyze the design choice of the VP module in TC-CLIP on the few-shot and zero-shot setting.
Vision token selection. We compare using the context tokens with using [CLS] or Global Average Pooled (GAP) tokens from every frame. Using context tokens yields the best result, demonstrating that the proper contextualization from the video is essential for transferring the video information to the text side.
Text prompting design. We observe that (a) naively adding learnable prompt vectors without video instance conditioning does not enhance the zero-shot transferability and decreases the average accuracy by 0.2%p. (b) In contrast, employing VP design with [CLS] tokens consistently improves the accuracy across all datasets, and (c) using context tokens further enhances the performance, resulting in a 1.6%p gain. We also compare VP with (d) vision-text late-fusion design, i.e., the cross-attention of context tokens and the final representation of the text embedding. This design shows worse performance in UCF-101 and K-600 datasets than our VP, demonstrating the superiority of our design choice.
5.4 Qualitative Analysis
Context token visualization. Fig. 6 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP. In this video, the informative regions regarding the action of “disc golfing” in each frame, including the person and the disc, are selected as seed tokens. To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region. Note that a single context token (highlighted in red) successfully tracks the disc across multiple frames.
Attention visualization. Fig. 7 visualizes the attention map of ViFi-CLIP [26] and our TC-CLIP on the SSv2 dataset. In this video where two hands grab objects and then move away, ViFi-CLIP [26] fails to attend to the hands from the middle of the sequence and misinterprets the action as “colliding with each other”. In contrast, our TC-CLIP considers the temporal context across the sequence by its design, and thus consistently attends to the hands throughout the entire video and correctly predicts the action as “moving away from each other”. See more visualizations in Appendix 0.E.

### v2

@@ 2.4 Video-conditional Prompting (VP)
Figure 5: Video-conditional Prompting (VP) module. Video information from the context tokens is injected into the text prompt vectors using a cross-attention mechanism, generating instance-level prompts that make up for the lack of textual semantics.
The Video-conditional Prompting (VP) module further leverages the comprehensive video information derived from the visual domain for text encoding. We apply a cross-attention between prompt vectors and context tokens to enrich the information in the prompt vectors as illustrated in Fig. 5. Let $\textbf{c}^{l-1}$ and $\textbf{p}^{l-1}$ be class name tokens and learnable prompt vectors from the $(l-1)^{\text{th}}$ layer of the text encoder, respectively. We derive temporally contextualized prompt vectors $\tilde{\mathbf{p}}^{l-1}$ by passing the layer-normalized prompt tokens and context tokens through a cross-attention layer and an FFN layer as follows:
@@
$\hat{\mathbf{p}}^{l-1}$ $=\text{MHCA}(\text{LN}_{p}(\mathbf{p}^{l-1}),\text{LN}_{s}(\mathbf{s}_{\text{proj}}^{l}))+\mathbf{p}^{l-1},$ (13)
$\tilde{\mathbf{p}}^{l-1}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{p}}^{l-1}))+\hat{\mathbf{p}}^{l-1},$ (14)
where $\text{SG}(\cdot)$ is a stop-gradient function, $\mathbf{W}_{\text{vis}}$ is a weight matrix of CLIP to linearly project vision representations onto a common vision-language latent space, and $\text{MHCA}(\cdot,\cdot)$ is a multi-head cross-attention operation for interactions across modalities, accepting text prompt vectors as queries and vision features as keys and values. The VP module $f_{\theta_{\text{VP}}}(\cdot,\cdot)$ is defined by a composition of Eq. (13) and Eq. (14), and executed before the last layer of text encoder $f_{\theta_{c}}$. Finally, the new formulation of our encoding process in the text modality is given by
@@ 2.5 Training Objective
TC-CLIP learns to maximize the similarity of video representations $\mathbf{v}$ and text representations $\mathbf{c}$ for matching pairs in a mini-batch via the cross-entropy loss as $\mathcal{L}=-\sum_{i}\log\frac{\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{j})/\tau)},$ where $\tau$ is a learnable temperature parameter. Our model is fully fine-tuned in an end-to-end manner.
@@ 3 Experiments
Table 2: Comparison with state-of-the-arts on zero-shot action recognition. All the models are trained on Kinetics-400 and directly evaluated on other datasets. WE indicates the weight-space ensemble between the fine-tuned model and CLIP, adopted for all applicable models for fair comparisons. † denotes results reproduced using our implementation. The best results are in bold-faced numbers, and the second-best ones are underlined. Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
Method WE HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
@@
ViFi-CLIP [35] 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 92.1 $\pm$ 0.3 67.3
TC-CLIP (Ours) 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 93.2 $\pm$ 0.2 68.9
ActionCLIP [41] 51.9 $\pm$ 0.5 74.2 $\pm$ 1.0 67.5 $\pm$ 1.2 90.7 $\pm$ 0.1 64.5
ViFi-CLIP [35] 52.2 $\pm$ 0.7 81.0 $\pm$ 0.9 73.9 $\pm$ 0.5 93.3 $\pm$ 0.3 69.0
Open-VCLIP [44] 53.9 $\pm$ 1.2 83.4 $\pm$ 1.2 73.0 $\pm$ 0.8 93.2 $\pm$ 0.1 70.1
TC-CLIP (Ours) 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 94.4 $\pm$ 0.2 71.0
@@
MAXI [25] 52.3 $\pm$ 0.7 78.2 $\pm$ 0.8 71.5 $\pm$ 0.8 92.5 $\pm$ 0.4 67.3
OST [5] 55.9 $\pm$ 1.2 79.7 $\pm$ 1.1 75.1 $\pm$ 0.6 94.6 $\pm$ 0.2 70.2
FROSTER [11] 54.8 $\pm$ 1.3 84.8 $\pm$ 1.1 74.8 $\pm$ 0.9 - 71.5
TC-CLIP (Ours) 56.0 $\pm$ 0.3 85.4 $\pm$ 0.8 78.1 $\pm$ 1.0 95.7 $\pm$ 0.3 73.2
@@
Table 4: Comparison with state-of-the-arts on base-to-novel generalization. All the models are directly fine-tuned from CLIP. † results are taken from [11].
@@
Table 5: Fully-supervised action recognition results on Kinetics-400. Views means (temporal clips) $\times$ (spatial crops), and F denotes number of frames.
Method Top-1 Top-5 F Views
ActionCLIP [41] 83.8 96.2 32 10 $\times$ 3
X-CLIP [30] 84.7 96.8 16 4 $\times$ 3
Vita-CLIP [43] 82.9 96.3 16 4 $\times$ 3
ViFi-CLIP [35] 83.9 96.3 16 4 $\times$ 3
OST [5] 83.2 - 16 1 $\times$ 1
TC-CLIP (Ours) 85.2 96.9 16 4 $\times$ 3
[…]
Method Zero Few B2N Full Params GFLOPs TP
ActionCLIP [41] 64.5 46.5 46.0 83.8 143.7M 567 20
X-CLIP [30] 60.6 50.2 49.5 84.7 169.7M 288 36
Vita-CLIP [43] 63.7 - - 82.9 161.8M 307 30
ViFi-CLIP [35] 69.0 52.9 55.5 83.9 124.3M 285 38
Open-VCLIP [44] 70.1 - 56.5 - 124.3M 308 29
TC-CLIP (Ours) 71.0 54.8 58.5 85.2 127.5M 304 24
TC-CLIP (Ours)§ 70.7 54.4 58.6 84.9 127.5M 291 34
We conduct experiments on 5 video benchmarks: Kinetics-400 [17] & 600 [2], HMDB-51 [22], UCF-101 [39], and Something-Something v2 (SSv2) [10]. Following [35], our evaluation protocols include zero-shot, few-shot, base-to-novel generalization, and fully-supervised action recognition tasks. We adopt CLIP with ViT-B/16 for all experiments and our baseline is ViFi-CLIP [35]. All models are trained using 4 NVIDIA Tesla V100 GPUs. More details are in the appendix.
3.1 Quantitative Comparison
We mainly compare our method with CLIP-based video recognition models: Vanilla CLIP [33], ActionCLIP [41], A5 [15], X-CLIP [30], Vita-CLIP [43], ViFi-CLIP [35], Open-VCLIP [44], OST [5], and FROSTER [11]. For the fair comparisons with approaches based on Large Language Model (LLM) with text augmentation [25, 5, 11], we produce two versions of our results: one using the original action category names (colored in blue) and the other adopting the LLM-rephrased category names obtained from FROSTER [11] (colored in purple). Note that experiments on the SSv2 dataset do not involve LLM-rephrasing.
Zero-shot action recognition. Table 2 exhibits the zero-shot generalization ability of several models, where they are trained on K-400 and then directly evaluated on individual datasets. For fair comparisons with recent models [44, 25, 5, 11], we employ weight-space ensembling (WE) for all applicable models except those freezing a backbone during fine-tuning. Specifically, the weights of both vision and text encoders are linearly interpolated between CLIP and the fine-tuned model as $\theta_{w}=(1-w)\cdot\theta_{\text{CLIP}}+w\cdot\theta_{\text{fine-tuned}}$. TC-CLIP consistently outperforms others across all datasets, showing its superior generalization ability.
Few-shot action recognition. We verify the learning capacity of our method under a challenging few-shot scenario. In Table 3, models are directly fine-tuned from CLIP on each dataset using $K$-shot samples, where $K$ is 2, 4, 8, and 16. TC-CLIP achieves the best performance with large margins from ViFi-CLIP [35].
Base-to-novel generalization. Similarly, models are directly fine-tuned from CLIP using the base classes of each dataset and evaluated for both base and novel classes. Table 4 reports top-1 accuracies on the base and novel classes with their harmonic mean (HM). TC-CLIP performs the best on the novel classes and HM across all datasets, especially showing solid results on the SSv2 dataset.
@@ 3.1 Quantitative Comparison
Fully-supervised action recognition. Table 5 shows performance comparison results under the fully-supervised setting, where the models are trained and evaluated both on the K-400 dataset. TC-CLIP achieves top-1 accuracy of 85.2% in the validation split, improving 1.3%p over our baseline ViFi-CLIP [35].
[…]
@@
Table 7: Component-wise ablations on the zero-shot setting. $\Delta$ denotes the average top-1 accuracy gain over baseline.
Without weight-space ensembling With weight-space ensembling
Case HMDB-51 UCF-101 K-600 All ($\Delta$) HMDB-51 UCF-101 K-600 All ($\Delta$)
Baseline 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 67.3 52.2 $\pm$ 0.7 81.0 $\pm$ 0.9 73.9 $\pm$ 0.5 69.0
(a) +TC 53.6 $\pm$ 0.2 78.6 $\pm$ 1.0 71.8 $\pm$ 0.7 68.0 (+0.7) 54.3 $\pm$ 0.6 81.9 $\pm$ 1.0 75.5 $\pm$ 1.0 70.6 (+1.6)
(b) +VP 53.2 $\pm$ 0.8 80.5 $\pm$ 0.7 71.6 $\pm$ 0.9 68.4 (+1.1) 53.4 $\pm$ 0.8 82.0 $\pm$ 0.9 74.7 $\pm$ 0.7 70.0 (+1.0)
(c) +TC+VP 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 68.9 (+1.6) 54.2 $\pm$ 1.1 82.9 $\pm$ 0.9 75.8 $\pm$ 0.4 71.0 (+2.0)
Table 8: Effect of TC with various token aggregation strategies. TC consistently outperforms the frame-wise attention baseline across several different token selection and merging methods. $K$-shot action recognition results are reported with the top-1 accuracy averaged over $K=2,4,8,16$. Default settings are marked in gray.
(a) Seed token selection strategy.
Case HMDB UCF SSv2 All ($\Delta$)
Baseline 62.6 89.2 8.7 53.5
No selection 62.8 89.8 9.7 54.1 (+0.6)
Head-wise key norm 62.3 89.8 9.8 54.0 (+0.5)
Averaged key norm 62.5 89.4 9.3 53.7 (+0.2)
Head-wise CLS attn. 63.4 89.9 9.7 54.3 (+0.8)
Averaged CLS attn. 63.4 90.2 9.9 54.5 (+1.0)
Patch saliency [6] 62.9 90.3 9.6 54.2 (+0.7)
ATS [9] 63.5 90.3 9.8 54.5 (+1.0)
(b) Context token summarization strategy.
Case HMDB UCF SSv2 All ($\Delta$)
Baseline 62.6 89.2 8.7 53.5
No merge 57.2 85.6 7.7 50.2 ($-3.3$)
Random merge 58.8 87.1 7.5 51.2 ($-2.3$)
K-means [26] 62.1 89.7 9.0 53.6 (+0.1)
DPC-KNN [14] 63.3 90.2 9.8 54.4 (+0.9)
Bipartite soft matching [16, 1] 63.4 90.2 9.9 54.5 (+1.0)
Bipartite w/ attention weights 62.9 89.8 9.9 54.2 (+0.7)
Bipartite w/ saliency weights [6] 62.4 89.9 9.6 54.0 (+0.5)
Table 9: TC design ablation. We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. Bias is defined in Eq. (8).
(a) Positional embedding design.
@@
Spatial embedding 62.9 90.0 9.8 54.2
Joint space-time embedding 63.2 90.2 9.8 54.4
Spatial embedding + Bias 63.4 90.2 9.9 54.5
Joint embedding + Bias 62.9 90.2 9.8 54.3
(b) Seed token ratio $\alpha$.
@@
Table 10: Text prompting design ablation on the zero-shot setting. All the models are evaluated without the weight ensemble.
@@
3.2 Analysis and Discussion
This section examines the design choices and impact of each component in our model: Temporal Contextualization (TC) and Video-conditional Prompting (VP). We mainly adopt the zero- and few-shot settings and report the average of top-1 accuracy with $K=2,4,8,16$ for the $K$-shot setup. In addition to the analyses discussed in this subsection, we present more analyses and qualitative results in the supplementary document.
Component-wise ablation. Table 7 shows the impact of TC and VP on our baseline in the zero-shot setting. Integrating TC gives an average gain of 0.7%p over the baseline and the gap increases to 1.6%p after adopting WE; WE is more favorable to our approach than the baseline. Adopting VP also leads to a substantial gain of 1.1%p, highlighting its own contribution. When both VP and TC are applied to the baseline, an average improvement goes up to 1.6%p, which finally leads to 2.0%p gain after applying WE.
Token aggregation strategies. In Table 8, we verify the effectiveness of TC across diverse token aggregation methods. Experiments are conducted on the few-shot setting using the baseline model with TC. (a) While TC still works well without token selection, we observe that collecting informative seed tokens based on token importance, such as attention or saliency scores, improves the quality of encoded tokens by suppressing the background. (b) Directly using the seed tokens without merging reduces performance due to the extrapolation issue. The degradation with random merging also highlights the requirement of token clustering based on relevance. Finally, consistent gains from various token merging approaches verify the robustness of TC regardless of algorithms.
Positional embedding. Table 9(a) shows that using the proposed learnable bias (Eq. (8)) with spatial positional embedding yields the best result. We conjecture that the bias effectively consolidates the local frame-level information and global video-level information in a layer-wise and head-wise manner.
Number of seed and context tokens. While TC is not sensitive to the choice of $\alpha$, as shown in Table 9(b), we picked $\alpha=0.3$ as our default value, i.e., using 30% of total tokens as seed tokens. In Table 9(c), the context token number $k$ is chosen to set a modest amount of merging degree.
Text prompting design. In Table 10, we observe that (a) a naïve integration of learnable prompt vectors without video instance conditioning is not particularly helpful for the zero-shot transferability, rather decreasing the average accuracy. In contrast, (b) employing VP design with [CLS] tokens consistently improves the accuracy across all datasets, and (c) using context tokens further enhances the performance, resulting in a 1.6%p gain. We also compare VP with (d) vision-text late-fusion design, i.e., the cross-attention of context tokens and the final representation of the text embedding. This design performs worse in UCF-101 and K-600 datasets than our VP, verifying the effectiveness of our design choice.